ELUCIDATION OF EXTRA-MECHANICAL EFFECTS IN REDUCTION OF INFRARED AND RAMAN SPECTRA OF SMALL MOLECULES IN GASEOUS SAMPLES

\underline{J. F. OGILVIE}, Centre for Experimental and Constructive Mathematics, Department of Mathematics, Simon Fraser University, 8888 University Drive, Burnaby, BC V5A 1S6 Canada.